# Estimating energy levels from lattice QCD correlation functions using a transfer matrix formalism

### v1

@@ Abstract
We present an efficient method for estimating energy levels from lattice QCD correlation functions by computing the eigenvalues of the transfer matrix of the corresponding lattice QCD Hamiltonian. This method is equivalent to the Lanczos procedure introduced recently [1], but more general and simpler to implement. We demonstrate the efficacy of the method by calculating the two lowest energy levels of a large number of hadrons including a few nuclei. While the signal-to-noise problem does not improve significantly, the extracted energy levels are found to be more reliable, compared to those extracted by conventional methods. Within a given statistical dataset the proposed method can effectively account for both statistical uncertainties and systematic errors, including those arising from the fitting window, and hence can reliably be employed in lattice QCD calculations.
preprint: TIFR/TH/24-24
@@ I Introduction
Extraction of energy levels from two-point correlation functions is an essential component in most lattice QCD studies. The methods involved in that need to be reliable and efficient not only for determining the energy spectra accurately but also for probing and understanding the internal structures of hadrons, particularly for any precision study. In general, by exploiting the asymptotic decay of two-point correlation functions, $C(\tau)\sim e^{-E_{n}\tau}$, one can extract the ground state, $E_{0}$, at a sufficiently large distances from its source of excitation ($\tau\rightarrow\infty$). Over the years, better methodologies for extracting energy levels have been developed within the framework of the generalized eigenvalue problem (GEVP) [2, 3]. These approaches use a large variational basis of interpolating fields to estimate the energy levels more reliably. All these methods have produced remarkable results, particularly for stable hadrons [4, 5], achieving per-mille accuracy in many cases and even predicting subatomic particles before their experimental discovery [6, 7, 8]. More recently, these methods have also been extensively applied to study resonance properties and scattering phenomena [9]. However, a significant challenge in these procedures remains the signal-to-noise-ratio (SNR) problem in Euclidean correlation functions, as pointed out by Parisi [10] and Lepage [11]. This issue, intrinsic to lattice QCD calculations, becomes more pronounced for systems involving multi-hadron states, and becomes a hindrance particularly for nuclei [12, 13, 14, 15]. Overcoming this bottleneck remains a key focus in improving the precision and reliability of lattice QCD studies.
@@
Recently, an alternative procedure has been introduced [1, 16] based on computation of eigenvalues of the transfer matrix. Since the transfer matrix is infinite dimensional, a recursive procedure using Lanczos method was employed, leveraging Kaniel-Paige-Saad (KPS) bound on energy estimation [17, 18]. The efficacy of the above method was shown using toy model calculations as well as for pion and nucleon. However, as noted there, a major challenge with this method is filtering out spurious eigenvalues, which can introduce subjectivity and systematic errors into the estimation of energy levels.
Following this development, here we propose a similar method to estimate the energy levels from the eigenvalues of the transfer matrix. Specifically, we employ another Krylov space-based method for eigenvalue extraction from two-point correlation functions of lattice QCD  [^1] During the final stage of drafting this article, another article was published on arXiv (Ref. [19]), which also utilized the similar method as in this work. However, we used a different filtering procedure of eigenvalues and the error estimation is also different.. Our method is straightforward to implement and can be seamlessly extended to the generalized eigenvalue problem (GEVP) [2, 3], offering enhanced simplicity and ease of implementation. We employ a variation of the Prony generalized eigenvalue method [20, 21, 22, 23], and we refer this method as Transfer matrix GEVP (TGEVP). We demonstrate the potency of the method by extracting the two lowest eigenvalues from the various two-point functions of a variety of hadrons, including nuclei. Similar to Ref. [19] we also find that the energy estimates obtained with this method are identical to those derived using the oblique Lanczos algorithm outlined in [1]. Below we describe the method.
@@ II Methodology
In lattice QCD calculations, a two-point correlation function, corresponding to an interpolating field $\mathcal{O}$, between a source ($\mathbf{x_{i}},t_{i}$) and sink ($\mathbf{x_{f}},t_{f}$) is given by,
@@
where $|\Omega\rangle$ represents the vacuum state of the theory. In terms of the Hamiltonian matrix ${\mathcal{H}}$, in general we also express a two-point function as,
$C(t)=\langle\Omega|\mathcal{O}(0)e^{-\hat{\mathcal{H}}t}\overline{\mathcal{O}}(0)|\Omega\rangle,$ (2)
If $\hat{\mathcal{T}}\equiv e^{-\hat{\mathcal{H}}}$ is the transfer matrix in Euclidean space-time, in terms of the states, $|\chi\rangle\equiv\overline{\mathcal{O}}(0)|\Omega\rangle$ created by the interpolating field, we can write the above expression as,
$C(t)=\langle\chi|\hat{\mathcal{T}}^{t}|\chi\rangle.$ (3)
In lattice QCD simulations, however, $C(t)$ is computed only at a finite number of discrete times, specifically at $t\in{a,2a,\cdots,n_{T}a}$, where $a$ is the lattice spacing and $n_{T}$ is the temporal extent of the lattice. This leads to
$C(na)=\langle\chi|\hat{\mathcal{T}}^{na}|\chi\rangle,\text{where$n\in{1,2,\cdots,n_{T}}$}.$ (4)
For simplicity, we assume the lattice units to be $a=1$, reducing the above equation to,
@@
The eigenvalues of the transfer matrix corresponds to the eigenstates of the Hamiltonian $\mathcal{H}$. Usually, while the transfer matrix is infinite dimensional, in classical computation one always needs to work with a truncated transfer matrix evaluated on a finite-dimensional basis. . Here we choose the vectors in Krylov-subspace, $\mathcal{K}_{m}$, as a set of linearly-independent basis vectors, which are given below,
@@
where $\lambda_{n}^{m}$’s and $x^{m}_{n}$’s are eigenvalues and eigenvectors of the transfer matrix truncated on the $(m+1)$ dimensional Krylov-subspace (we provide the proof in the Appendix A). Note that, $V^{m}$ and $T^{m}$ both are positive-definite matrices for a given source-sink setup. Hence the eigenvalues, $\lambda_{n}^{m}$, in Eq. (9) are real and positive. The energies of the eigenstates of the truncated transfer matrix are given by, $E^{m}_{n}=-\log\lambda_{n}^{m}$. Since the eigenenergies are positive, we expect $0\leq\lambda_{n}^{m}\leq 1$  [^2] As, described in [1], thermal states appear in the spectrum of $\mathcal{T}^{m}$ with $\lambda_{n}^{m}>1$. However we are not interested in these states here. .
@@
In the absence of statistical noise in a correlation function, $C(n)$, selecting eigenvalues across different iterations of the matrix $T^{m}$ is straightforward. The eigenvalues computed in the $m$-th iteration will be arranged in increasing order, $E^{m}_{0}<E^{m}_{1}<E^{m}_{2}<\cdots<E^{m}_{m}\;$, where $E^{m}_{0}$ represents the ground-state energy estimate for the $m$-th iteration, while $E^{m}_{i}$ for $1\leq i\leq m$ corresponds to the subsequent excited states. In lattice QCD simulations, however, the correlation functions are evaluated stochastically using the Monte Carlo path integral approach with finite statistics. This introduces statistical noise, causing fluctuations in the computed matrices $C^{m}$ and $T^{m}$ which can then deviate from their positive-definite structure, resulting spurious eigenvalues in the spectrum of the $T^{m}$. This issue is particularly problematic in systems with a dense spectrum, as the spurious eigenvalues can disrupt the proper ordering of energy levels, making the extraction of reliable energy levels challenging. This issue has already been pointed out in Ref. [1], where a filtering procedure with Cullum-Willoughby method [24] was employed to eliminate these spurious eigenvalues. However, as noted there, such filtering procedure provides only a partial solution to the problem which can make the eigenvalue estimation, particularly their errors, unreliable.
II.1 Procedure for spurious eigenvalue filtration
To address the issue of spurious eigenvalues, we employ a pivoting technique that iteratively refines the eigenvalue histogram by weighting high-confidence peaks based on Gaussian filtering and full-width half-maximum criteria. Below, we outline the steps involved in the eigenvalue filtration process:

### v2

@@ Abstract
We present an efficient method for extracting energy levels from lattice QCD correlation functions by computing the eigenvalues of the transfer matrix associated with the lattice QCD Hamiltonian. While mathematically and numerically equivalent to the recently introduced Lanczos procedure [1], our approach introduces a novel prescription for removing spurious eigenvalues using a kernel density estimator (KDE) and Gaussian-convoluted histogram method. This strategy yields a robust and stable estimate of the energy spectrum, outperforming the Cullum-Willoughby filtering technique in efficiency. In addition, we detail how this method can be applied to extract overlap factors from two-point correlation functions, as well as matrix elements from three-point functions with a current insertion. Furthermore, we extend the methodology to accommodate correlation matrices constructed from a variational basis of operators, with its Block formulation. We demonstrate the efficacy of this framework by computing the two lowest energy levels for a broad range of hadrons, including several nuclei. Although the signal-to-noise ratio is not significantly improved, the extracted energy levels are found to be more reliable than those obtained with conventional techniques. Within a given statistical ensemble, the proposed method effectively captures both statistical uncertainties and systematic errors, including those arising from the choice of fitting window, making it a robust and practical tool for lattice QCD analysis.
@@ I Introduction
Extraction of energy levels from two-point correlation functions is an essential component in most lattice QCD studies. The methods involved in that process need to be reliable and efficient not only for determining the energy spectra accurately but also for probing and understanding the internal structures of hadrons, particularly for any precision study. In general, by exploiting the asymptotic decay of two-point correlation functions, $C(\tau)\sim e^{-E_{n}\tau}$, one can extract the lowest-lying energy level $E_{0}$ that couples to a given interpolating operator at sufficiently large temporal separations from the source ($\tau\rightarrow\infty$). Over the years, better methodologies for extracting energy levels have been developed within the framework of the generalized eigenvalue problem (GEVP) [2, 3]. These approaches use a large variational basis of interpolating fields to estimate the energy levels more reliably. All these methods have produced remarkable results, particularly for stable hadrons [4, 5], achieving per-mille accuracy in many cases and even predicting subatomic particles before their experimental discovery [6, 7, 8]. More recently, these methods have also been extensively applied to study resonance properties and scattering phenomena [9]. However, a significant challenge in these procedures remains the signal-to-noise-ratio (SNR) problem in Euclidean correlation functions, as pointed out by Parisi [10] and Lepage [11]. This issue, intrinsic to lattice QCD calculations, becomes more pronounced for systems involving multi-hadron states, and becomes a hindrance particularly for nuclei [12, 13, 14, 15]. Overcoming this bottleneck remains a key focus in improving the precision and reliability of lattice QCD studies.
@@
Recently, a novel procedure was introduced [1, 16] for extracting energy levels by computing eigenvalues of the transfer matrix. Given the infinite-dimensional nature of this operator, a recursive Lanczos method was employed, making use of the Kaniel-Paige-Saad (KPS) bound for energy estimation [17, 18]. The efficacy of the above method was shown using toy model calculations as well as for pion and nucleon. However, as noted there, a major challenge with this method is filtering out spurious eigenvalues, which can introduce subjectivity and systematic errors into the estimation of energy levels.
Following this developments, we propose a similar approach to extract energy levels from the eigenvalues of the transfer matrix in lattice QCD. Specifically, we introduce a Krylov subspace-based method for eigenvalue extraction from two-point correlation functions. We show that the oblique Lanczos method introduced in Ref. [1] can be reformulated as a generalized eigenvalue problem (GEVP), as a variant of the Prony GEVP method [19, 20, 21, 22, 23, 24]. We henceforth refer to this framework as the Transfer Matrix GEVP (TGEVP). Both analytically and numerically, we demonstrate the equivalence of this formulation with the oblique Lanczos approach. The same conclusion has also been reached in Refs. [25, 26] [^1] Ref. [25] appeared simultaneously with this work, while Ref. [26] came subsequently..
A major challenge in these TGEVP/Lanczos-type methods lie in identifying and removing spurious eigenvalues, which—if not address properly—can bias the extracted energy levels and increase systematic uncertainty. Ref. [1] advocated a bootstrap generalization of the Cullum–Willoughby (CW) test for filtering these spurious modes. However, as noted there, the CW test can be imperfect in the presence of noise. Additionally, we find its performance depends critically on tuning three hyperparameters that vary across datasets. This makes it difficult to apply uniformly without dataset-specific calibration. As we demonstrate in Subsec. V.2, even when tuned, the CW method may fail to produce stable estimates.
To address this, we propose a data-driven, non-parametric filtering method based on estimating the statistical distribution of eigenvalues using Gaussian-smoothed histograms and adaptive kernel density estimators (KDEs). Peaks in the resulting distribution and their full-width at half-maxima (FWHM) are utilized to identify and exclude outlier eigenvalues that deviate significantly from the dominant physical modes. While histogram- and KDE-based peak estimation and distribution estimations are well-established in other fields (see e.g., [27, 28, 29, 30, 31]), and also utilized in lattice QCD in different contexts [32, 33], here we employ such techniques in spectral filtering of spurious eigenvalues of TGEVP matrix. Our histogram/KDE-based filtration method for extracting physical eigenvalues from noisy TGEVP spectra provides a data-driven, hyperparameter-light alternative to traditional filtering techniques (such as the CW test), and achieves comparable or superior stability across a variety of datasets without the need for manual tuning, as demonstrated in Sec. V with the numerical results.
The remainder of the draft is structured as follows. In Sec. II, we present the details of the proposed TGEVP method. Sec. III outlines how histogram and kernel density estimation (KDE) can be employed to estimate the distribution of eigenvalues, following established techniques from the statistics literature [34]. In Sec. IV, we provided a flowchart to describe the filtration workflow used to identify and remove spurious eigenvalues. Results obtained using TGEVP with the proposed filtering procedure are presented in Sec. V. In Subsec. V.1, we compare the outcomes with those from the standard effective mass approach and Prony’s method. A detailed comparison between the Lanczos and TGEVP approaches is provided in Subsec. V.2. Subsecs. V.3V.5 present results for a wide range of hadrons, including light and heavy baryons and nuclei, for a variety of lattice ensembles, illustrating the broad applicability and effectiveness of our method for extracting energy levels. In Subsec. V.6, we examine correlations across TGEVP iterations for testing the reliability of the method. In Subsec. V.7, we discuss the extraction of overlap factors, along with supporting numerical results, for studying decay properties of hadrons. Subsec. V.8 demonstrates how the TGEVP framework can be incorporated into multi-operator GEVP methods to access excited states and finite-volume amplitude analysis. Finally, Sec. VI provides a summary of our findings along with concluding remarks and future directions. The mathematical equivalence between the TGEVP and oblique Lanczos approaches is shown in Appendix A. The asymptotic convergence properties of histogram and KDE-based distribution estimation are discussed in Appendix B. Additional results and supporting figures are presented in Appendix C. In Appendix D we discuss the impact of complex eigenvalues in estimating the physical energy levels.
@@ II Methodology
In lattice QCD calculations, the two-point correlation function for an interpolating field $\mathcal{O}$, between a source at $(\mathbf{x}_{i},t_{i})$ and a sink at $(\mathbf{x}_{f},t_{f})$, is given by
@@
where $|\Omega\rangle$ denotes the vacuum state of the theory. In Euclidean space-time, this can also be expressed in terms of the Hamiltonian $\mathcal{H}$ as
$C(t)=\langle\Omega|\mathcal{O}(0)\,e^{-\hat{\mathcal{H}}t}\,\overline{\mathcal{O}}(0)|\Omega\rangle.$ (2)
Defining the Euclidean transfer matrix as $\hat{\mathcal{T}}=e^{-a\hat{\mathcal{H}}}$, and the state $|\chi\rangle\equiv\overline{\mathcal{O}}(0)|\Omega\rangle$, we obtain
$C(t)=\langle\chi|\hat{\mathcal{T}}^{t/a}|\chi\rangle.$ (3)
In lattice simulations, time is discretized as $t=na$, where $n\in\{1,2,\ldots,n_{T}\}$, and the above simplifies to
@@
The eigenvalues of the transfer matrix correspond to the eigenstates of the Hamiltonian $\mathcal{H}$. In order to study the transfer matrix explicitly, one must work with a truncated version defined on a finite-dimensional basis. This is in contrast to path integral Monte Carlo methods, which sample configurations and can implicitly encode contributions from the full, effectively infinite-dimensional Hilbert space. In our implementation, we choose the Krylov subspace, $\mathcal{K}_{m}$, as a set of linearly independent basis vectors for this purpose of truncating $\hat{\mathcal{T}}$. These basis vectors are outlined below.
@@
where $\lambda_{n}^{m}$’s and $x^{m}_{n}$’s are eigenvalues and eigenvectors of the transfer matrix truncated on the $(m+1)$ dimensional Krylov-subspace. We provide the proof in the Appendix A. Note that, $V^{m}$ and $T^{m}$ both are positive-definite matrices for a given source-sink setup. Hence the eigenvalues, $\lambda_{n}^{m}$, in Eq. (8) are real and positive. The energies of the eigenstates of the truncated transfer matrix are given by, $E^{m}_{n}=-\log\lambda_{n}^{m}$. Since the eigenenergies are positive, we expect $0\leq\lambda_{n}^{m}\leq 1$  [^2] As, described in [1], thermal states appear in the spectrum of $\mathcal{T}^{m}$ with $\lambda_{n}^{m}>1$. However we are not interested in those states here, and so are removed.. Note in this formalism, changing the problem as a standard GEVP enables the use of well-established linear algebra routines, avoiding the need for a recursive oblique Lanczos implementation. Moreover, as we will show later that overlap factors and matrix elements can directly be extracted from the GEVP eigenvectors, unlike in the oblique Lanczos method, which requires an additional basis transformation. This makes TGEVP conceptually simpler and easier to implement.
@@
In the absence of statistical noise in a correlation function, $C(n)$, selecting eigenvalues across different iterations of the matrix $T^{m}$ is straightforward. The eigenvalues computed in the $m$-th iteration will be arranged in increasing order, $E^{m}_{0}<E^{m}_{1}<E^{m}_{2}<\cdots<E^{m}_{m}\;$, where $E^{m}_{0}$ represents the ground-state energy estimate for the $m$-th iteration, while $E^{m}_{i}$ for $1\leq i\leq m$ corresponds to the subsequent excited states. In lattice QCD simulations, however, the correlation functions are evaluated stochastically using the Monte Carlo path integral approach with finite statistics. This introduces statistical noise, causing fluctuations in the computed matrices $C^{m}$ and $T^{m}$ which can then deviate from their positive-definite structure, resulting spurious eigenvalues in the spectrum of the $T^{m}$. This issue is particularly problematic in systems with a dense spectrum, as spurious eigenvalues can disrupt the proper ordering of energy levels, making the extraction of reliable results challenging. The problem was also highlighted in Ref. [1], where a filtering procedure based on the Cullum–Willoughby (CW) method [35] was employed to remove such spurious modes. However, as noted in that work, the CW test can be unreliable in the presence of statistical noise. Moreover, we find that its effectiveness depends sensitively on the tuning of three hyperparameters, which vary significantly across different datasets. This sensitivity makes it difficult to apply the method in a consistent and fully automated way.
To overcome these limitations, we introduce a filtration strategy based on Gaussian-convoluted histograms and adaptive kernel density estimators (KDEs). These methods are data-driven, and while the former requires minimal manual tuning, KDE is fully automated offering a more robust and broadly applicable alternative. Using these methods, from a lattice two-point function, one can extract the energy levels reliably without subjective tuning of hyperparameters. A full technical description of our filtration methodology is presented in the following section, and subsequently we demonstrate the efficacy and usefulness of these methods over any other methods.
III Filtering Spurious Eigenvalues in GEVP
Our strategy for removing spurious eigenvalues arising in the solution of the generalized eigenvalue problem (GEVP) is based on the observation that such modes typically manifest as outliers relative to the central bulk of the eigenvalue distribution of the transfer matrix under bootstrap resampling. If left unfiltered, these spurious modes can significantly bias both the extracted central values and the associated uncertainties in an statistical analysis.
To address this issue, we identify the dominant peaks in the eigenvalue distribution and define peak-specific thresholds to eliminate outliers lying beyond the main support of each peak. A key ingredient in this procedure is a reliable estimation of the eigenvalue density obtained from the bootstrap-resampled correlators. To this end, we utilize two complementary non-parametric techniques: Gaussian-convoluted histograms and adaptive kernel density estimators (KDEs). These provide smooth and robust estimates of the underlying eigenvalue distribution without assuming a specific parametric form.
From the estimated eigenvalue density, we identify the prominent peaks corresponding to the dominant eigenvalue modes. The FWHM of each peak is measured, yielding an estimate of the spread of the eigenvalue distribution around the central modes. This FWHM serves as a quantitative measure of the uncertainty or fluctuation of the eigenvalue modes in the correlation function. Using the FWHM estimates, we then apply a cut (to be discussed later) to the data by discarding any eigenvalues that fall outside the expected range, defined as a factor of the FWHM from the center of the peak. This cut removes outliers, eigenvalues that lie far from the main distribution and could distort further analysis, and hence yields a cleaner eigenvalue dataset, enhancing the reliability of subsequent computations.
The use of histograms and kernel density estimators as non-parametric tools for distribution estimation from stochastic samples is well-established in statistics. For detailed treatments, we refer to Refs. [36, 37, 34, 38]. These techniques are also utilized in various other scientific areas for peak detection and peak-width estimation in noisy distributions [27, 28, 29, 30, 31]. To the best of our knowledge, however, this work constitutes the first systematic application of such methods to outlier removal in the spectral analysis of lattice QCD correlators.
III.1 Gaussian-Convoluted Histogram
As a first step toward estimating the eigenvalue density from bootstrap-resampled data, we begin with the construction of a histogram—one of the simplest non-parametric estimators of a probability density function (PDF). Given a set of eigenvalues $\{x_{1},x_{2},\cdots,x_{n}\}$, the histogram groups the data into bins and normalizes the counts to approximate the underlying distribution. The resulting estimator $\hat{f}_{h}(x)$, with bin width $h$ and origin $x_{0}$, is defined by:
$\hat{f}_{h}(x)$ $=$ $\frac{1}{nh}\sum_{i=1}^{n}\mathcal{I}(x_{i}\in B_{j}),$ (9)
where $\text{where }B_{j}=[x_{0}+(j-1)h,x_{0}+jh),$
where $\mathcal{I}(\cdot)$ denotes the indicator function. Under suitable conditions, this estimator converges to the true density as $n\to\infty$ and $h\to 0$ [34, 38]. We summarize its statistical properties, including bias, variance, and convergence rate, in Appendix B. Note that, in this case the error in $\hat{f}_{h}(x)$ scales as $n^{-2/3}$ [34], which is slower than that expected from a more refined estimator.
While intuitive and easy to implement, the histogram suffers from several known drawbacks. It is piecewise constant by construction, sensitive to the choice of bin origin, and prone to introducing artificial discontinuities. These limitations become especially problematic in our context, where statistical fluctuations are present, and fine spectral structure must be preserved without over-interpreting noise.
To address these issues, we smooth the histogram by convolving it with a Gaussian kernel, producing a Gaussian-convoluted histogram. This approach retains the non-parametric nature of the histogram while mitigating its roughness. The smoothed estimator $\tilde{f}_{h}(x)$ is given by,
[…]
where the Gaussian kernel takes the form,
$K_{\sigma}(x)=\frac{1}{\sqrt{2\pi}\sigma}\exp\left(-\frac{x^{2}}{2\sigma^{2}}\right),$ (11)
and $\sigma$ is the kernel width that controls the degree of smoothing. This convolution serves to suppress high-frequency noise and resolve genuine features of the distribution by replacing abrupt bin edges with a continuous smooth curve. In particular, it improves the reliability of peak detection, which is central to our filtration strategy.
However, as with all smoothing techniques, an appropriate care must be taken in choosing the parameter $\sigma$. If $\sigma$ is too small, the estimator remains noisy and prone to false peaks; if too large, it can over-smooth the distribution, merging nearby peaks or shifting modal locations. We therefore explore a range of $\sigma$ values and assess the stability of key features, such as peak positions and full width at half maximum (FWHM), to guide the selection of both $\sigma$ and the histogram bin width $h$.
Importantly, since our application to correlation functions uses these peaks only to define outlier-rejection thresholds, the robustness of the final mean and percentile estimates remains intact, and this is valid even if the smoothed FWHM marginally overestimates the true peak width. This makes the Gaussian-convoluted histogram a practical and stable tool in the first stage of our density estimation pipeline.
III.2 Adaptive Kernel density Estimator
In addition to the Gaussian-convoluted histogram method, we employ kernel density estimation (KDE) to obtain a smooth, non-parametric estimate of the eigenvalue density. As we show below this method is found to be completely data-driven and more robust for the purpose of extracting energy levels from two-point correlation functions. By constructing the density as a superposition of smooth kernels centered at each data point, KDE avoids manual tuning of $\sigma$ associated with histogramming, such as bin widths and edge placements. We explain the method below. Given a bootstrap-resampled set of eigenvalues, $\{x_{1},x_{2},\cdots,x_{n}\}$, the kernel density estimate $\hat{f}_{h}(x)$ is given by,
$\hat{f}_{h}(x)=\frac{1}{nh}\sum_{i=1}^{n}K\left(\frac{x-x_{i}}{h}\right),$ (12)
where $K(\cdot)$ is a continuous kernel function and $h$ is the bandwidth that governs the degree of smoothing. For our purposes, we adopt the Gaussian kernel,
$K(u)=\frac{1}{\sqrt{2\pi}}e^{-u^{2}/2},$ (13)
yielding the Gaussian KDE,
$\hat{f}_{h}(x)=\frac{1}{nh\sqrt{2\pi}}\sum_{i=1}^{n}\exp\left(-\frac{(x-x_{i})^{2}}{2h^{2}}\right).$ (14)
This approach has several advantages. It produces a smooth, differentiable estimate of the density that is less sensitive to local noise and discretization effects than histograms. Moreover, it enables accurate identification of distinct peaks and their widths, both of which are essential for our automated outlier rejection procedure. Importantly, KDE does not require the manual specification of bin edges or origins, thereby removing a key source of arbitrariness in traditional histogram-based approaches. We discuss its statistical properties, including bias, variance, and convergence rate, in Appendix B. As discussed there, another advantage of KDE over histogram is its faster asymptotic convergence as $n^{-4/5}$ [36, 34, 38] with sample size $n$. It can be shown that, under weak assumptions, there is no non-parametric estimator that converges at a faster rate than that of the kernel estimator [39].
Just as with histograms, however, the effectiveness of KDE depends crucially on the choice of smoothing parameter $h$. Too small a bandwidth results in noisy estimates dominated by statistical fluctuations, while too large a bandwidth merges neighboring modes and washes out genuine features. We therefore adopt a data-driven strategy to determine the optimal smoothing scale, especially suited to the potentially multimodal structure of bootstrap-resampled eigenvalue spectra. We explain the implementation procedure below in details:
Bandwidth Selection and Local Adaptivity
To balance resolution with stability, we employ an adaptive KDE scheme based on the balloon estimator [37, 40]. Unlike fixed-bandwidth KDE, which imposes a uniform smoothing scale across the entire spectrum, this approach modulates the bandwidth locally, guided by the density of the data around each evaluation point. Such adaptivity is crucial in our context, where the spectral landscape may feature narrow peaks associated with low-lying energy states alongside broader, and also flatter structures characteristic of excited levels. By allowing the degree of smoothing to vary with local data density, the estimator captures fine structure without over-fitting statistical noise in the tails. Below we elaborate the procedure in a few sequential steps.